HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ON THE SPECTROSCOPY OF ACYLIUM- AND THIOACYLIUM IONS

\underline{S. THORWIRTH} , O. ASVANY, T. SALOMON, M. BAST,

P. C. SCHMID, S. SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany

Acylium- and thioacylium ions, R-CO^{+} and R-CS^{+}, are a class of molecular ions of relevance for astrochemistry that have received relatively little attention from molecular spectroscopy so far. Triggered by our first spectroscopic study of the polyatomic HC_{3}O^{+} and HC_{3}S^{+} species,^{1} it was not long until these two ions as well as the methyl variant CH_{3}CO^{+} were detected in space.^{2} Here, we would like to present a status report on the spectroscopic studies of acylium- and thioacylium ions at Cologne that make use of various action spectroscopy schemes in combination with state-of-the-art ion trap machinery.

1doi:10.1080/00268976.2020.1776409S. Thorwirth, M. E. Harding, O. Asvany, S. Brünken, P.

Jusko, K. L. K. Lee, T. Salomon, M. C. McCarthy, and S. Schlemmer, Descendant of the X-ogen carrier and a ”mass of 69”: Infrared action spectroscopic detection of HC_{3}O^{+} and HC_{3}S^{+}, Mol. Phys.

118, e1776409 (2020).

2doi:10.1051/0004-6361/202040076J. Cernicharo, C. Cabezas, S. Bailleux, L. Margulès, R.

Motiyenko, L. Zou, Y. Endo, C. Bermúdez, M. Agúndez, N. Marcelino, B. Lefloch, B. Tercero, and P. de Vicente, Discovery of the acetyl cation, CH_{3}CO^{+}, in space and in the laboratory, Astron. Astrophys. 46, L7 (2021) and references therein.

p-number: p149  Submitted on Wed, 14 Jun 2023 19:18:09 +0200